### v1

# Invisible neutron decay and light BSM particles

Abstract
In Standard Model Effective Field Theory (SMEFT) invisible neutron decay arises from $d=12$ operators. Adding new, light particles to the field content of the SM, such as right-handed neutrinos, allows to construct operators for invisible neutron decay at much lower dimensions. Observing invisible neutron decay, if nucleon decays with charged leptons remain absent, would therefore point towards the existence of new neutral degrees of freedom. Here, we discuss four cases: (i) Adding right-handed neutrinos to the SM; (ii) a right-handed neutrino and an axion-like particle; (iii) a right-handed neutrino and a (nearly) massless singlet scalar; and (iv) a right-handed neutrino and a light $Z^{\prime}$. We give the general tree-level decomposition for the resulting d=(79)d=(7-9) operators for invisible neutron decay. We also briefly discuss LHC searches related to the exotic states found in these UV completions.
Nucleon decays, SMEFT, light sterile neutrinos, LHC
I Introduction
Searches for nucleon decays by Super-Kamiokande have pushed limits to $10^{34}$ yr and above for several decay channels. The best known examples are probably $p\to\ell^{+}\pi^{0}$ ($\ell=e,\mu$) Super-Kamiokande:2020wjk and $n\to\bar{\nu}K^{0}$ Super-Kamiokande:2025ibz , which are motivated by Grand Unified Theories (GUTs). Less known are decays to three charged leptons, where Super-Kamiokande also gives limits in excess of $10^{34}$ yr now Super-Kamiokande:2020tor . Limits on many other channels, however, are less stringent, sometimes by orders of magnitude, see Takhistov:2016eqm ; ParticleDataGroup:2024cfk . On the other hand, nucleon decay searches form an important part of the agenda of next generation underground detectors such as JUNO JUNO:2015zny , Hyper-Kamiokande Hyper-Kamiokande:2018ofw or DUNE DUNE:2020lwj and significant improvements in sensitivity on many nucleon decay modes are expected from these experiments in the foreseeable future.
Traditionally, searches for proton decay were motivated by the predictions from GUTs, such as the classical $SU(5)$ Georgi:1974sy , for a review see for example Nath:2006ut . More recently, however, many theoretical papers have taken a more agnostic view and studied nucleon decays from the point of view of effective field theory (EFT) deGouvea:2014lva ; Kobach:2016ami ; Hambye:2017qix ; Fonseca:2018ehk ; Helo:2018bgb ; Helo:2019yqp ; Heeck:2019kgr ; Beneito:2023xbk ; Gargalionis:2024nij ; Heeck:2024jei ; Li:2024liy ; Li:2025slp ; IBeneito:2025nby ; Heeck:2025btc ; Heeck:2025uwh . [^1] Here we also cite some pioneering works on nucleon decays in SMEFT Weinberg:1979sa ; Wilczek:1979hc ; Weinberg:1980bf ; Weldon:1980gi ; Abbott:1980zj ; Claudson:1981gh and the studies on the decompositions of the relevant effective operators Bowes:1996xy ; Kovalenko:2002eh ; Arnold:2012sd ; Assad:2017iib ; deBlas:2017xtg ; Li:2023cwy . In Standard Model Effective Field Theory (SMEFT), one can classify different decay modes according to the dimensionality of the corresponding operators. For example, two-body decays with $\Delta(B+L)=2$ are described by $d=6$ operators, $\Delta(B-L)=2$ are $d=7$, genuine three-body decays need at least $d=9$ operators Heeck:2019kgr .
In the current paper, we are interested in invisible neutron decay. Considering only SM particles, the only possibility for a neutron to decay into “nothing” is the channel $n\to 3\nu$. [^2] Here both $\nu$ and $\bar{\nu}$ are mentioned as $\nu$. SMEFT operators with three leptons have been discussed in Hambye:2017qix . In SMEFT, the lowest dimensional operator with three leptons has $d=9$, one example is ${\cal O}\propto(QL)({\bar{L}}{\bar{L}})(d_{R}d_{R})$. Ref. Hambye:2017qix lists a total of 16 operators of this type at $d=9$. Some of these $d=9$ operators can generate the decay $n\to 3\nu$. However, all operators that can do so, always also generate decays to charged leptons with very similar rates. Since nucleon decay searches are much more sensitive to final states with charged leptons Super-Kamiokande:2020tor than to $n\to 3\nu$ ParticleDataGroup:2024cfk , $d=9$ operators can not generate invisible neutron decay at observable rates.
In SMEFT one has to go up to $d=12$, to find for the first time an operator capable of generating invisible neutron decay, without the corresponding decays to charged leptons, as already realized by Weinberg more than 40 years ago Weinberg:1980bf . Weinberg gave as an example the operator Weinberg:1980bf :
$\mathscr{L}^{d=12}_{\rm W}=\frac{c_{\text{W}}}{\Lambda^{8}}\epsilon_{IJK}\epsilon^{ij}\epsilon^{kl}\epsilon^{mn}\left(\overline{u_{R}}^{I}L_{i}\right)\left(\overline{d_{R}}^{J}L_{k}\right)\left(\overline{d_{R}}^{K}L_{m}\right)H_{j}H_{l}H_{n},$ (1)
where $\Lambda$ is the typical scale of the origin of this effective operator and $c_{\text{W}}$ is the Wilson coefficient. In Eq. (1) we have suppressed generation indices for simplicity, $\epsilon_{IJK}$ and $\epsilon^{ij}$ are anti-symmetric tensors for the contraction of $SU(3)_{c}$ and $SU(2)_{L}$ indices. This operator leads to the decay mode $n\to 3\nu$. For this $d=12$ operator one can estimate the nucleon decay width for the 3-body final state roughly as Heeck:2019kgr
$\Gamma\sim\frac{\beta_{h}^{2}m_{N}^{5}\langle H^{0}\rangle^{6}}{6144\pi^{3}\Lambda^{16}}$ $\Rightarrow$ $\frac{1}{9\cdot 10^{29}\;{\rm yr}}\Big(\frac{13.4\;{\rm TeV}}{\Lambda}\Big)^{16}.$ (2)
Here, $\beta_{h}\simeq 0.014$ GeV3 is the nuclear matrix element Aoki:2017puj , $m_{N}$ the nucleon mass, $\langle H^{0}\rangle\simeq 174$ GeV the vacuum expectation value of the SM Higgs field, and $c_{\text{W}}$ is taken to be unity.
For invisible neutron decay the particle data booklet ParticleDataGroup:2024cfk cites an old Kamiokande result (from 1993) Kamiokande:1993ivj ($4.9\cdot 10^{26}$ yr), but there are more recent and stronger bounds set by SNO SNO:2003lol ($2\cdot 10^{29}$), KamLAND KamLAND:2005pen ($5.8\cdot 10^{29}$ yr, cf. also Kamyshkov:2002wp ) and two analysis by the SNO+ collaboration SNO:2018ydj ($2.5\cdot 10^{29}$ yr) and SNO:2022trz ($9.0\cdot 10^{29}$ yr). For the nearer and farther future sensitivity estimates have been given by JUNO JUNO:2024pur ($5\cdot 10^{31}$ yr) and for THEIA (80kt), that could ultimately reach $\sim 4\cdot 10^{32}$ yr according to their proposal Theia:2019non .
However, there exists a simple possibility to generate invisible neutron decay with lower dimensional operators: We can extend the SM particle content with new light degrees of freedom. In this paper we discuss four different possibilities. First we add (i) (at least two) right-handed neutrinos $N_{R}$, also sometimes called sterile neutrinos in the literature, cf. Acero:2022wqg ; Dasgupta:2021ies ; Diaz:2019fwt ; Boser:2019rta ; Dentler:2018sju ; Kopp:2013vaa ; Abazajian:2012ys for reviews and current bounds, motivated from various neutrino experiments LSND:2001aii ; MiniBooNE:2010idf ; MiniBooNE:2020pnu ; Elliott:2023cvh ; Barinov:2021asz ; Dentler:2017tkw ; STEREO:2022nzk . In addition, the introduction of such light degrees of freedom is often discussed in the context of cosmology, e.g., the stringent bound on neutrino masses from recent cosmological observations DESI:2024mwx ; DESI:2024hhd ; DESI:2025zgx ; DESI:2025ejh maybe be alleviated by the mixing between active neutrinos and massless sterile neutrinos Farzan:2015pca ; Escudero:2022gez ; Benso:2024qrg ; Ota:2024byu . In $N_{R}$SMEFT, for a list of basis operators see Bhattacharya:2015vja ; Liao:2016qyd ; Li:2021tsq , invisible neutron decay operators appear at $d=9$.
Second, we add (ii) a right-handed neutrino and an axion-like particle (ALP), $a$. The axion is one of the best-motivated light new physics fields Weinberg:1977ma ; Wilczek:1977pj . It arises as a consequence of the breaking of the Peccei-Quinn symmetry introduced to solve the strong CP problem Peccei:1977hh ; Peccei:1977ur . For experimental searches and phenomenology of ALPs, cf. e.g., Graham:2015ouw ; Brivio:2017ije ; Bauer:2017ris ; Irastorza:2018dyq ; Bauer:2018uxu ; Gavela:2019wzg ; Merlo:2019anv ; Alda:2025uwo and for cosmological aspects of ALPs, cf. e.g., Cadamuro:2011fd ; Arias:2012az ; Marsh:2015xka ; Dror:2021nyr ; Chadha-Day:2021szb ; Adams:2022pbo . The basis operators of aaSMEFT, the effective theory with the SM particle content plus an ALP, are listed in Song:2023lxf ; Song:2023jqm ; Grojean:2023tsd . In $aN_{R}$SMEFT, i.e. SMEFT with $N_{R}$s and ALPs, invisible neutron decay operators appear at d=8d=8.333Basis operators with a given particle content can be constructed up to a given mass dimension $d$ by AutoEFT Harlander:2023ozs or also Sym2Int Fonseca:2017lem
[…]
In this paper, we will give the operators for invisible neutron decay for all four cases, discussed above, estimate the neutron decay half-lives and give the complete list of the possible UV completions for these operators at tree-level. For the different BSM particles found in these UV models we will also briefly discuss current LHC limits.
This introduction would not be complete without mentioning that invisible neutron decay has been discussed in some earlier papers. For example, in Dvali:1999hn the authors speculated about “decays to nothing” in models with extra time dimensions. The authors of Mohapatra:2002ug considered universal extra dimensions and an extended gauge group, to have the neutron decay dominantly into $n\to\nu\bar{\nu}_{s}\bar{\nu}_{s}$, where $\nu_{s}$ is some sterile neutrino. The effective operators relevant to $n\rightarrow\bar{\nu}_{s}\bar{\nu}^{\prime}_{s}\bar{\nu}^{\prime\prime}_{s}$ are listed in Girmohanta:2019fsx in the context of an extra dimension model. In Girmohanta:2020eav invisible neutron decay was studied in a left-right symmetric model with large extra dimensions. Some recent works have considered also the possibility of a “dark” (invisible) neutron decay Fornal:2018eol ; Barducci:2018rlx ; Cline:2018ami ; Berezhiani:2018udo ; Fornal:2020gto in relation to the neutron life-time anomaly Czarnecki:2018okw . And, finally, a particular UV model for invisible di-neutron decay to two singlet fermions has been studied in  Hao:2022xyj , where the authors discussed also possible constraints from neutron stars.
The rest of this paper is organized as follows. In section 2 we will introduce the different model variants in a bit more detail, discuss some generalities, define the operators and estimate the neutron decay half-lives for the different cases. In section 3 we will then decompose the operators to find the possible UV models at tree-level. Section 4 discusses a number of different LHC searches and how they can be used or reinterpreted to derive limits on the BSM states found in section 3. We then close with a short discussion and summary. Some tables for the decompositions will be relegated to the appendix for the sake of keeping the main text more compact.
II Setup: Models and effective operators
In this section we will describe the theoretical setup. We consider four different extensions of the SM particle content. For each case, we briefly describe the model. We then define the operators relevant for invisible neutron decay. We give an estimate of the neutron decay half-live in each case. Finally, in an additional subsection we discuss a possible caveat. The decomposition of the operators, defined here, is discussed in section 3.
II.1 Variant I: Right-handed neutrinos
In the standard model neutrinos are massless. While there are many extensions of the SM that can explain neutrino masses, as observed in oscilllation experiments, maybe the simplest possibility is to add right-handed neutrinos to the SM particle content.
Massive neutrinos can be either Dirac or Majorana particles. In the classical type-I seesaw model Minkowski:1977sc ; Yanagida:1979as ; Mohapatra:1979ia ; Schechter:1980gr right-handed neutrinos have a Majorana mass term and the Lagrangian can be written as:
$\mathscr{L}_{Y}=(Y_{\nu}^{*})_{\alpha j}\overline{N_{Rj}}L_{\alpha}\cdot H+\frac{1}{2}(M_{M})_{jj}\overline{{N_{Rj}}^{c}}N_{Rj}+\quad{\rm h.c.}$ (3)
Note that $M_{M}$ can be taken diagonal without loss of generality. Majorana neutrinos, on the other hand, violate lepton number by two units. Whether such a $\Delta L=2$ term is allowed or not, will depend on the ultra-violet completion of the operators defined below. We note, however, that the invisible neutron decay $n\to 3\bar{N}$ violates $L$ by three units, thus, Majorana masses need not to be allowed in general and neutrinos could be Dirac particles, as far as invisible neutron decay is concerned.
For the fit to current neutrino data, at least two copies of $N_{R}$ are needed and both Dirac and Majorana neutrinos can explain existing data. We will consider three copies of $N_{R}$, to mirror the three SM generations, but doing an exact fit to oscillation data is irrelevant for the neutron decay we are interested in. The only condition for the decay to occur is that the right-handed neutrino masses are smaller than $m_{n}/3$.
Two operators can be constructed at level $d=9$, that contribute to invisible neutron decay:
${\cal O}_{1}^{N}$ $=$ $\frac{c_{\alpha\beta\gamma ijk}^{(1)}}{\Lambda^{5}}\left(\overline{{u_{R\alpha}}^{c}}d_{R\beta}\right)\left(\overline{{d_{R\gamma}}^{c}}N_{Ri}\right)\left(\overline{{N_{Rj}}^{c}}N_{Rk}\right),$ (4)
${\cal O}_{2}^{N}$ $=$ $\frac{c_{\alpha\beta\gamma ijk}^{(2)}}{\Lambda^{5}}\left(\overline{{Q_{\alpha}}^{c}}Q_{\beta}\right)\left(\overline{{d_{R\gamma}}^{c}}N_{Ri}\right)\left(\overline{{N_{Rj}}^{c}}N_{Rk}\right).$ (5)
Here, $\alpha,\beta,\gamma$ are SM generation indices, while $i,j,k$ runs over the generation of $N_{R}$. To keep the expressions simpler, we have suppressed $SU(3)_{c}$ (and $SU(2)_{L}$) indices. Here, and in all cases discussed below the colour of the three quarks must be contracted with a completely anti-symmetric tensor, $\epsilon_{IJK}$. It is straightforward to show that both of these operators vanish in case there is only one copy of $N_{R}$, we have checked this also with Sym2Int Fonseca:2017lem .
Assuming the mass of the right-handed neutrinos to be negligible and one of the $c_{111ijk}^{(1/2)}=1$, we can estimate the partial half-life for the invisible neutron decay as
$\Gamma(n\rightarrow\bar{N}_{i}\bar{N}_{j}\bar{N}_{k})\simeq\frac{\beta_{h}^{2}m_{n}^{5}}{6144\pi^{3}\Lambda^{10}}\simeq\frac{1}{9\cdot 10^{29}\text{yr}}\left(\frac{180\text{TeV}}{\Lambda}\right)^{10}.$ (6)
Here again, $\beta_{h}\simeq 0.014$ GeV3 is the nuclear matrix element Aoki:2017puj and the half-live is normalized to the current bound from SNO+ SNO:2022trz . We can apply Eq. (6) to the different existing and projected half-life limits and obtain estimates for the sensitivity to the new physics scale $\Lambda$:
$\Gamma\sim$ $\begin{cases}\displaystyle\frac{1}{4.9\cdot 10^{26}\text{yr}}\left(\frac{84\text{TeV}}{\Lambda}\right)^{10}\text{ Kamiokande (PDG)},\\ \displaystyle\frac{1}{5.0\cdot 10^{31}\text{yr}}\left(\frac{270\text{TeV}}{\Lambda}\right)^{10}\text{ JUNO (future)},\\ \displaystyle\frac{1}{4.0\cdot 10^{32}\text{yr}}\left(\frac{330\text{TeV}}{\Lambda}\right)^{10}\text{ THEIA (future)}.\end{cases}$ (7)
II.2 Variant II: A right-handed neutrino and a ALP
Axion-like particles (ALPs) appear in many extensions of the SM. While the ALP mass is a free parameter, the couplings of the ALPs with SM fields are protected by an approximate classical shift symmetry, as is the case for the classical axion. Many different studies of properties of and searches for ALPs have been published, see for example Brivio:2017ije ; Bauer:2021mvw ; Bauer:2017ris ; Biekotter:2025fll . Here we only briefly mention that ALPs could also explan the dark matter Jaeckel:2010ni ; Alekhin:2015byh and/or be long-lived particles Alekhin:2015byh ; Alimena:2019zri .
The ALP Lagrangian up to $d=5$ contains the following terms Georgi:1986df
$\mathscr{L}_{a}$ $=$ $\frac{1}{2}\partial_{\mu}a\partial^{\mu}a-\frac{1}{2}m_{a}^{2}a^{2}-\sum_{X}\frac{c_{X\widetilde{X}a}}{\Lambda}aX^{\mu\nu}\widetilde{X}_{\mu\nu}-\sum_{\psi}\frac{c_{\psi a}}{\Lambda}\partial_{\mu}a({\overline{\psi}}\gamma^{\mu}\psi).$ (8)
Here $X^{\mu\nu}$ stands for any of the field strength tensors of the SM, i.e. $X=B,W,G$, and $\widetilde{X}_{\mu\nu}$ is its dual. $m_{a}$ is the ALP mass, in principle a free parameter. Note, however, that one usually assumes the ALP to be the pseudo-Goldstones of a spontaneously broken global symmetry, thus $m_{a}$ can be naturally small compared to the scale of symmetry breaking. $\psi$ are the SM fermions and for the case we are interested in also $\psi=N_{R}$. It is important to note that the ALP couples only derivatively to fermions.
There are two $d=8$ $B$-violating operators containing a right-handed singlet fermion, $N_{R}$, and an ALP, $a$:
${\cal O}_{1}^{a}$ $=$ $\frac{c_{\alpha\beta\gamma}^{(1,a)}}{\Lambda^{4}}(\partial_{\mu}a)\left(\overline{N_{R}}\gamma^{\mu}d_{R\alpha}\right)\left(\overline{{u_{R\beta}}^{c}}d_{R\gamma}\right),$ (9)
${\cal O}_{2}^{a}$ $=$ $\frac{c_{\alpha\beta\gamma}^{(2,a)}}{\Lambda^{4}}(\partial_{\mu}a)\left(\overline{N_{R}}\gamma^{\mu}d_{R\alpha}\right)\left(\overline{{Q_{\beta}}^{c}}Q_{\gamma}\right).$
One copy of $N_{R}$ is enough to form these operators, thus for simplicity we have not added a generation index to $N_{R}$. However, two $N_{R}$ would still be needed, if one wants to explain neutrino data.
There are two more $d=8$ $B$-violating operators, containing an ALP, that can be written down with only SM fields:
${\cal O}_{3}^{a}$ $=$ $\frac{c_{\alpha\beta\gamma\delta}^{(3,a)}}{\Lambda^{4}}(\partial_{\mu}a)\left(\overline{L_{\alpha}}d_{R\beta}\right)\left(\overline{{Q_{\gamma}}^{c}}\gamma^{\mu}d_{R\delta}\right),$ (10)
${\cal O}_{4}^{a}$ $=$ $\frac{c_{\alpha\beta\gamma\delta}^{(4,a)}}{\Lambda^{4}}(\partial_{\mu}a)\left(\overline{e_{R\alpha}}\gamma^{\mu}d_{R\beta}\right)\left(\overline{{d_{R\gamma}}^{c}}d_{R\delta}\right).$
The operators in Eq. (9) can generate invisible neutron decays, without accompanying charged lepton final states, while ${\cal O}_{3}^{a}$ will always also have charged leptons in the nucleon decays. We have listed ${\cal O}_{4}^{a}$ for completeness, it does not give a contribution to invisible neutron decay.
The operators in Eqs. (4) and (5) violate $(B+L)=4$. Thus, it is easy to argue that the standard $d=6$ 2-body $(B+L)=2$ decays are absent on symmetry grounds for the invisible neutron decay to three $N_{R}$’s. However, the situation is different for the case of an ALP. ALPs do not usually carry either $B$ or $L$, as can be seen from Eq. (8). For an ALP with $(B,L)=(0,0)$, however, ${\cal O}_{1,2}^{a}$ both have (BL)=2(B-L)=2. One can write down operators at $d=7$ with (BL)=2(B-L)=2, such as, for example, ${\cal O}\propto(\overline{N_{R}}\gamma^{\mu}d_{R\alpha})(\overline{{Q_{\beta}}^{c}}(iD_{\mu})Q_{\gamma})$. Since this type of operators leads to 2-body decays, such as $n\to\pi^{0}+E\hskip-6.00006pt/\;\>$ and $p\to\pi^{+}+E\hskip-6.00006pt/\;\>$, one would expect the invisible neutron decay to be a very sub-dominant decay mode in this case.
A possible way out of this conclusion is based on the following argument. If we assign $a$ the quantum numbers (B,L)=(1,1)(B,L)=(-1,1), ${\cal O}_{1,2}^{a}$ both conserve $(B+L)$ and (BL)(B-L), thus none of the (BL)=2(B-L)=2 operators need to be present in the theory. For this to be possible, however, the “standard” $d=5$ ALP interaction terms in Eq. (8) need to be absent. It is possible to write down UV completions that fulfil these assignments.
For the operators ${\cal O}_{1,2}^{a}$, the width for the two-body decay $n\rightarrow Na$ can be estimated to be roughly
$\Gamma(n\rightarrow Na)\simeq\frac{\beta_{h}^{2}m_{n}^{3}}{32\pi\Lambda^{8}}\simeq\frac{1}{9.0\cdot 10^{29}\text{yr}}\left(\frac{9.6\cdot 10^{6}\text{GeV}}{\Lambda}\right)^{8}.$ (11)
The much larger scale found here, compared to the case $n\rightarrow 3\bar{N}$, see Eq. (6), makes it unlikely that the UV completions for these operators can ever be tested in accelerator experiments.
II.3 Variant III: A right-handed neutrino and a light scalar
The third possibility for generating operators for the invisible neutron decay with BSM fields is to add a right-handed neutrino and a light scalar to the SM. We discuss this case only briefly, since we believe it to be less motivated and add it only for completeness of the discussion.
For a model with a light singlet scalar, $\phi$, and right-handed neutrinos there are two $d=7$ operators contributing to invisible neutron decay:
${\cal O}_{1}^{\phi}$ $=$ $\frac{c_{\alpha\beta\gamma}^{(1,\phi)}}{\Lambda^{3}}\phi\left(\overline{{N_{R}}^{c}}d_{R\alpha}\right)\left(\overline{{u_{R\beta}}^{c}}d_{R\gamma}\right),$ (12)
${\cal O}_{2}^{\phi}$ $=$ $\frac{c_{\alpha\beta\gamma}^{(2,\phi)}}{\Lambda^{3}}\phi\left(\overline{{N_{R}}^{c}}d_{R\alpha}\right)\left(\overline{{Q_{\beta}}^{c}}Q_{\gamma}\right).$
Similar to the ALP case, discussed above, if $\phi$ has $(B,L)=(0,0)$, the $d=6$ operators that one can construct from Eq. (12), eliminating $\phi$, would lead to dominant 2-body decays, rendering invisible neutron decay uninteresting. Asigning non-zero baryon and lepton number to $\phi$ allows to eliminate this problem technically.
The width $n\to\bar{N}\phi$ can be estimated to be roughly
$\Gamma(n\rightarrow\bar{N}\phi)\simeq\frac{\beta_{h}^{2}m_{n}}{32\pi\Lambda^{6}}\simeq\frac{1}{9.0\cdot 10^{29}\text{yr}}\left(\frac{2.1\cdot 10^{9}\text{GeV}}{\Lambda}\right)^{6}.$ (13)
II.4 Variant IV: A right-handed neutrino and a light vector
Finally, we briefly mention the possibility where a neutron decays to a right-handed neutrino and a SM-singlet vector boson $Z^{\prime}$, which is prompted by the following $d=7$ effective operators:
${\cal O}_{1}^{Z^{\prime}}$ $=$ $\frac{c_{\alpha\beta\gamma}^{(1,Z^{\prime})}}{\Lambda^{3}}Z^{\prime}_{\mu}\left(\overline{N_{R}}\gamma^{\mu}d_{R\alpha}\right)\left(\overline{{u_{R\beta}}^{c}}d_{R\gamma}\right),$ (14)
${\cal O}_{2}^{Z^{\prime}}$ $=$ $\frac{c_{\alpha\beta\gamma}^{(2,Z^{\prime})}}{\Lambda^{3}}Z^{\prime}_{\mu}\left(\overline{N_{R}}\gamma^{\mu}d_{R\alpha}\right)\left(\overline{{Q_{\beta}}^{c}}Q_{\gamma}\right).$
To avoid the possible $d=6$ nucleon decays and make the $d=7$ invisible neutron decay dominant, we need to assign charges (which may be the $B$ and $L$ numbers) to $Z^{\prime}$. The argument is the same as for the ALP case, cf. Sec. II.2.
Assuming $Z^{\prime}$ is a massless gauge boson under a new gauge symmetry, the decay rate is calculated to be
$\Gamma(n\rightarrow NZ^{\prime})\simeq$ $\frac{\beta_{h}^{2}m_{n}}{16\pi\Lambda^{6}}\simeq\frac{1}{9.0\cdot 10^{29}\text{yr}}\left(\frac{2.3\cdot 10^{9}\text{GeV}}{\Lambda}\right)^{6}.$ (15)
When $Z^{\prime}$ aquires a mass through the spontaneous breaking of some new gauge symmetry, the decay rate changes to:
$\Gamma\simeq\frac{\beta_{h}^{2}m_{n}^{3}}{32\pi\Lambda^{6}M_{Z^{\prime}}^{2}}\left(1-\frac{M_{Z^{\prime}}^{2}}{m_{n}^{2}}\right)^{2}\left(1+\frac{2M_{Z^{\prime}}^{2}}{m_{n}^{2}}\right).$ (16)
This equation seems to diverge in the limit of $M_{Z^{\prime}}\rightarrow 0$. However, for a gauge boson such divergencies are known to be unphysical. In a full-fledged model for the massive $Z^{\prime}$, one would need to specify also the symmetry breaking sector of the theory. The full calculation then includes the correct treatment of the would-be Goldstone boson of the theory and in the limit of $M_{Z^{\prime}}\ll m_{n}$ Eq. (16) will reduce to the same expression as for the scalar case Eq. (13), up to a coefficient. However, constructing such a complete model for the $Z^{\prime}$ is beyond the scope of the current paper.
II.5 A caveat
There is one important caveat to the above discussion: For any operator that annihilates three quarks to invisible particles it is possible to replace an initial state quark for a final state anti-quark. An example is shown in Fig. 1. Thus, four-body final states with a (neutral or charged) pion will always accompany three-body invisible decays. - And similarly for the two-body final states $N_{R}+a$ and $N_{R}+S$, three-body final states with one additional pion will be generated. [^5] Since one can replace an initial $d$-quark by a final state $\overline{s}$, producing a final state kaon, this allows, in principle, also sensitivity to 2nd generation indices in the operators.
No search for such three- and four-body final states exist. However, Super-Kamiokande has a search for $n\to\pi^{0}\bar{\nu}$ and $p\to\pi^{+}\bar{\nu}$. The current bounds are Super-Kamiokande:2013rwg :
$\Gamma(p\rightarrow\pi^{+}\bar{\nu})>$ $3.9\cdot 10^{32}\text{yr},$ (17)
$\Gamma(n\rightarrow\pi^{0}\bar{\nu})>$ $1.1\cdot 10^{33}\text{yr}.$ (18)
These are two-body decays and thus the pion momentum is fixed at roughly $p(\pi)\sim 460$ MeV. Super-Kamiokande uses this constraint in setting the limits. For the 4-body decays, on the other hand, the pion momentum takes a distribution, which depends in addition on the mass of the final state particles. Nevertheless, Super-Kamiokande provides sufficient information such that we can make a rough estimate of the limit for the 4-body decay $n\to\pi^{0}+3\bar{N}$.
[…]
Figure 1: $p(n)\rightarrow\pi^{+}(\pi^{0})+\text{missing}$ induced by the $d=9$ $u_{R}d_{R}d_{R}N_{R}N_{R}N_{R}$ operator.
For the four-body decays of Fig. 1, we can estimate the partial half-live, using results from Heeck:2019kgr , with appropriate replacements:
$\Gamma(n\rightarrow\pi^{0}\bar{N}_{i}\bar{N}_{j}\bar{N}_{k})\simeq$ $\frac{m_{n}^{7}W_{0}(\pi)^{2}}{737280\pi^{5}\Lambda^{10}}\simeq\frac{1}{3.3\cdot 10^{32}\text{yr}}\left(\frac{245\text{TeV}}{\Lambda}\right)^{10}$ (19)
Here the matrix element is taken to be $\left|W_{0}(\pi)\right|=\left|\langle\pi^{0}|(ud)_{R}d_{R}|n\rangle\right|=0.134\text{GeV}^{2}$ Aoki:2017puj . And again, we assume the Wilson coefficient is equal to one.
Comparing this estimate with with Eq. (7), we can see that the sensitivity of this pion mode to the new physics scale $\Lambda$ is actually better than the sensitivity of the invisible mode at SNO+ (180 TeV), and only moderately lower than the future sensitivity estimated for JUNO (270 TeV). Given these numbers, one must conclude that searches for $n\to\pi^{0}+E\hskip-6.00006pt/\;\>$ provide important constraints on searches for invisible neutron decay. Therefore, we encourage experimentalists to study the $n\rightarrow\pi^{0}+$“multiple missing” mode (independently from $n\rightarrow\pi^{0}\bar{\nu}$) and provide the bound. We stress again, that this is not only true for the case of the $d=9$ operator discussed here, but for all operators generating invisible neutron decay.
III High-energy completions
In the previous section we have defined the effective operators for invisible neutron decay for the four possible different model variants. In this section we will discuss how these operators can be generated in the UV, after integrating out heavy particles. Our aim here is to find systematically all possible UV models via the diagrammatic method. Since, however, the basic procedure is the same for all decompositions, we will discuss here only one decomposition of the operator ${\cal O}_{1}^{N}\propto u_{R}d_{R}d_{R}N_{R}N_{R}N_{R}$ in detail. Full tables for all other operators and decompositions are given in the appendix.
Figure 2: Topologies for tree-level realizations of the $d=9$ neutron decay operators, see Eqs. (4) and (5). Depending on the chiralities of the outer fermion fields, the dashed lines can be either a scalar, $S$, or vector, $V$. The mediator $F$ in Topology B must be introduced as a vector-like fermion, but when it is a SM singlet, it can also be a Majorana fermion.
For 6-fermion operators there are only two possible topologies at tree-level Bonnet:2012kh , using only renormalizable interactions. They are shown in Fig. 2. Note that here the arrows indicate the flow of the particle number. Internal particles can be either scalars, fermions or vectors. For Topology A there are exactly three possibilities to distribute the fermions of ${\cal O}_{1}^{N}$ to the outer legs of the diagram. We list these in Tab. 1.
Decomposition $S_{1}$ $S_{2}$ $S_{3}$ Comment Eq.
$(u_{R}d_{R})(d_{R}N_{Ri})(N_{Rj}N_{Rk})$ $(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{1}}},+1/3)$ $(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},-1/3)$ $(\boldsymbol{{\mathrm{1}}},\boldsymbol{{\mathrm{1}}},0)$ $S_{2}\neq S_{1}^{\dagger}$ to avoid $d=6$ (20)
$(u_{R}N_{Ri})(d_{R}d_{R})(N_{Rj}N_{Rk})$ $(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},+2/3)$ $(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{1}}},-2/3)$ $(\boldsymbol{{\mathrm{1}}},\boldsymbol{{\mathrm{1}}},0)$ $d_{R}d_{R}S_{2}^{\dagger}=0$
$(u_{R}N_{Ri})(d_{R}N_{Rj})(d_{R}N_{Rk})$ $(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},+2/3)$ $(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},-1/3)$ $(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},-1/3)$ $S_{3}\neq S_{2}$ to avoid $S_{1}S_{2}S_{2}=0$ (21)
Table 1: Decompositions for the operator $u_{R}d_{R}d_{R}N_{Ri}N_{Rj}N_{Rk}$ allowed with Topology A. The fields in a parenthesis form an interaction with the corresponding mediator field, $S_{1,2,3}$. For the explicit form of the interaction Lagrangians, cf. Eqs. (20) and (21).
For ${\cal O}_{1}^{N}$ and Topology A, all mediator fields are Lorentz scalars (indicated with symbol $S$). The charges of them under the SM gauge symmetries, in the order $(SU(3)_{c},SU(2)_{L},U(1)_{Y})$, are uniquely determined as given in the table. The interaction Lagrangians for the different cases can be written as:
$\mathscr{L}_{\text{A-1}}=$ $y_{ud}\epsilon^{IJK}(\overline{{u_{R}}^{c}})_{I\dot{a}}(d_{R})_{J}^{\dot{a}}(S_{1}^{\dagger})_{K}+y_{dN}(\overline{{d_{R}}^{c}})_{I\dot{a}}(N_{Ri})^{\dot{a}}(S_{2}^{\dagger})^{I}+y_{NN}(\overline{{N_{Rj}}^{c}})_{\dot{a}}(N_{Rk})^{\dot{a}}S_{3}^{\dagger}$
$+\mu(S_{1})^{I}(S_{2})_{I}S_{3},$ (20)
$\mathscr{L}_{\text{A-3}}=$ $y_{uN}(\overline{{u_{R}}^{c}})_{I\dot{a}}(N_{Ri})^{\dot{a}}(S_{1}^{\dagger})^{I}+y_{dNS_{2}}(\overline{{d_{R}}^{c}})_{I\dot{a}}(N_{Rj})^{\dot{a}}(S_{2}^{\dagger})^{I}+y_{dNS_{3}}(\overline{{d_{R}}^{c}})_{I\dot{a}}(N_{Rk})^{\dot{a}}(S_{3}^{\dagger})^{I}$
$+\mu\epsilon^{IJK}(S_{1})_{I}(S_{2})_{J}(S_{3})_{K},$ (21)
where the indices $I,J,K$ in the lower (upper) position are for $\boldsymbol{{\mathrm{3}}}$ ($\overline{\boldsymbol{{\mathrm{3}}}}$) in $SU(3)_{c}$, and $\dot{a}$ is for a right-handed 2-spinor. For case A-2 the interaction $d_{R}d_{R}S_{2}^{\dagger}$ vanishes identically, thus it is not a valid decomposition for ${\cal O}_{1}^{N}$. [^6] However, the interaction $d_{R}s_{R}S_{2}^{\dagger}$ does not vanish. Since operators are usually given in the flavour basis, while in the neutron the quarks are in the mass eigenstate basis, decomposition A-2 can contribute to the invisible neutron decay rate. The rate, however, will be suppressed by a factor $\sin^{2}\theta_{C}$, where $\theta_{C}$ is the Cabibbo angle. We will not discuss this possibility in further details. For a more compact notation, we have suppressed generation indices in these Lagrangians. The different Yukawas are to be understood as matrices of either dimensions ($3,3$), ($3,n$) or ($n,n$), where $n$ is the number of copies of right-handed neutrinos.
A few more comments are in order. In case A-1, described by the Lagrangian Eq. (20), the diquark, $S_{1}$, has the same SM charges as the leptoquark, $S_{2}^{\dagger}$. However, these two fields have to be distinct states. If they were identical, one could construct from $S_{1}$ a $d=6$ effective operator ${\cal O}\propto u_{R}d_{R}d_{R}N_{R}$. This operator generates the the 2-body decays $n(p)\rightarrow\pi^{0}(\pi^{+})\bar{N}$ with a significantly larger rate than that of the invisible neutron decay. To make $n\rightarrow 3\bar{N}$ the dominant decay mode, one must avoid all $d=6$ nucleon decays, i.e. $S_{1}$ and $S_{2}$ must be two different fields. This can be realized by assigning for case A-1 all $S_{i}$ definite baryon and lepton numbers. With the asignments $S_{1}=(2/3,0)$, $S_{2}=(1/3,1)$ and $S_{3}=(0,2)$, where the brackets stand for ($B,L$), all Yukawa couplings conserve baryon and lepton number trivially. The only $B$ and $L$ violating coupling would then be the soft parameter $\mu$. Since $\mu$ is a soft parameter, one can actually argue that it can be technically small in the sense of ’t Hooft’s naturalness, since in the limit of $\mu\to 0$ baryon and lepton numbers are conserved. A small value of $\mu$ would change the expected mass scale for the $S_{i}$ drastically. Instead of Eq. (6), for topology-A one can express the decay width as:
$\Gamma(n\rightarrow\bar{N}_{i}\bar{N}_{j}\bar{N}_{k})\simeq\frac{\beta_{h}^{2}m_{n}^{5}\mu^{2}}{6144\pi^{3}\Lambda^{12}}\simeq\frac{1}{9\cdot 10^{29}\text{yr}}\left(\frac{1\text{keV}}{\mu}\right)^{2}\left(\frac{2.4\text{TeV}}{\Lambda}\right)^{10}.$ (22)
Here, it is assumed that $m_{S_{1}}\simeq m_{S_{2}}\simeq m_{S_{3}}\simeq\Lambda$. Putting $\mu=\Lambda$, Eq. (22) reduces to Eq. (6), of course. However, this argument shows that it might not be completely hopeless to search for the BSM particles that appear in the decomposition of a $d=9$ operator at the LHC or the future FCC-hh.
The third case, case A-3, is similar to case A-1. However, see Eq. (21), case A-3 contains three leptoquarks, and two of them, $S_{2}$ and $S_{3}$, interact with $d_{R}$ and $N_{R}$. They therefore have the same SM charges. However, also in this case they must be two different fields, otherwise the triple scalar interaction vanishes. In this case, a simple lepton number assignment, as discussed for case A-1, is not sufficient to make $S_{2}$ and $S_{3}$ distinct. This problem can, on the other hand, be cured by adding a discrete symmetry, for example $\mathbb{Z}_{3}$, with the charge $\omega\equiv\text{e}^{\text{i}2\pi/3}$. We can assign $\omega$ to $N_{Rj}$ and $S_{2}$ and $\omega^{2}$ to $N_{Rk}$ and $S_{3}$. Under this assignment both, the triple scalar interaction as well as the Yukawas, are allowed under the discrete symmetry. However, the $\mu$-term will still break lepton and baryon numbers softly, thus guaranteeing that the dangerous 2-body nucleon decay modes are absent.
In the appendix, we list all mediator fields that appear in the high-energy completions of the $u_{R}d_{R}d_{R}N_{R}N_{R}N_{R}$ operator with Topology B in Fig. 2. There we also give the decompositions of the $QQd_{R}N_{R}N_{R}N_{R}$ operator Eq. (5) with both the topologies. We note that also for topology B it is, of course, possible to lower the expected mass scale. In EFT one usually estimates the scale $\Lambda$ for a Wilson coefficient $c\simeq 1$. However, for a $d=9$ operator $c\propto Y^{4}$, where $Y$ stands symbolically for any Yukawa coupling, see Fig. 2. Thus, even moderatly small Yukawa coouplings, will change the expections given in Eq. (6) drastically:
$\Gamma(n\rightarrow\bar{N}_{i}\bar{N}_{j}\bar{N}_{k})\simeq\frac{\beta_{h}^{2}Y^{8}m_{n}^{5}}{6144\pi^{3}\Lambda^{10}}\simeq\frac{1}{9\cdot 10^{29}\text{yr}}\left(\frac{Y}{0.01}\right)^{8}\left(\frac{4.5\text{TeV}}{\Lambda}\right)^{10}.$ (23)
In addition, the appendix contains tables of the decompositions for the effective operators in the other variants, which were presented in Secs. II.2-II.4. Similar to cases A-1 and A-3 discussed here for ${\cal O}_{1}^{N}$, in all cases one has to make sure, that the dangerous 2-body nucleon decays are absent by the use of some symmetry, otherwise the rate for the invisible neutron decay will be negligible.
$n\rightarrow 3\bar{N}$ $n\rightarrow Na(NZ^{\prime})$ $n\rightarrow\bar{N}\phi$
Mediator uudNNNuudNNN QQdNNNQQdNNN $\partial audd\bar{N}$ $(Z^{\prime}udd\bar{N})$ $\partial aQQd\bar{N}$ $(Z^{\prime}QQd\bar{N})$ $uudN\phi$ $QQdN\phi$
$S(\boldsymbol{{\mathrm{1}}},\boldsymbol{{\mathrm{1}}},0)$ $\checkmark$ $\checkmark$
$S(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},-1/3)$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
$S(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},+2/3)$ $\checkmark$ $\checkmark$ $\checkmark$
$S(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{2}}},+1/6)$ $\checkmark$
$V(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},-1/3)$ $\checkmark$ $\checkmark$
$V(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},+2/3)$ $\checkmark$
$V(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{2}}},+1/6)$ $\checkmark$ $\checkmark$ $\checkmark$
$F(\boldsymbol{{\mathrm{1}}},\boldsymbol{{\mathrm{1}}},0)$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
$F(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},-1/3)$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
$F(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},+2/3)$ $\checkmark$ $\checkmark$ $\checkmark$
$F(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{2}}},+1/6)$ $\checkmark$ $\checkmark$ $\checkmark$
Table 2: List of the mediators that appear in the decompositions of the effective operators relevant for the invisible neutron decay.
Table 2 lists all mediator fields that appear in the decompositions of the d=(79)d=(7-9) operators. There is a total of 11 fields, 4 scalars, 4 fermions and 3 vectors. Some fields appear in all operators and in all 4 model variants ($S_{3,1,-1/3}$, $F_{1,1,0}$ and $F_{3,1,-1/3}$), while some others appear only for a particular model variant ($S_{1,1,0}$, $V_{3,1,-1/3}$) or even only one particular operator ($S_{3,2,1/6}$ and $V_{3,1,2/3}$). Note that the fields appearing in the decompositions of the effective operator with a $Z^{\prime}$ are the same as the ones for the operators with an ALP, the corresponding column applies to both fields.
IV LHC phenomenology
In this section, we will briefly discuss a variety of LHC searches, that can be re-interpreted as limits for fields listed in Tab. 2. Naive expectations for the scale, $\Lambda$, of the d=(79)d=(7-9) operators, discussed in secton II, make it look quite unlikely that the LHC can find any positive signals for these states. However, as discussed around Eqs. (22) and (23), the estimates derived using EFT in section II might vastly overestimate the masses of the BSM fields in a full-fledged UV model. We therefore felt motivated to at least briefly discuss the current experimental status.
[…]
First, we mention that for the singlet fields, there is no production mode at the LHC other than decays of the coloured BSM states. Moreover, $F_{1,1,0}$ and $S_{1,1,0}$ will decay to $N_{R}$’s and thus be invisible, if they are lighter than the coloured states. Thus, there are no constraints on these two fields from LHC searches.
Let us then concentrate on the coloured fermions. Here, we follow largely the discussion in Cepedello:2024qmq . All three coloured fermions can have interactions with SM quarks and the Higgs:
${\cal L}_{\rm Y}$ $=$ $Y_{U_{R}}\overline{Q}F_{3,1,2/3}H^{\dagger}+Y_{D_{R}}\,\overline{Q}F_{3,1,-1/3}H+Y_{Q_{u}}\,\overline{F_{3,2,1/6}}u_{R}H^{\dagger}+Y_{Q_{d}}\,\overline{F_{3,2,1/6}}d_{R}H+{\rm h.c.}$ (24)
In Eq. (24) we have suppressed generation indices. The fermions $F_{3,1,-1/3}$, $F_{3,1,2/3}$ and $F_{3,2,1/6}$ must also have vector-like mass terms. After electro-weak symmetry breaking, Eq. (24) leads to mixing between heavy and SM quarks. This mixing will dominate their decay widths, unless the corresponding Yukawa couplings are tiny. If, on the other hand, some Yukawa couplings are large, this mixing may dominate also VLQ production at the LHC, see Fig. 3, left. In addition, VLQs can always be pair-produced, see Fig. 3 to the right, independent of the size of the mixing. Note that in this figure we show the VLQs decaying to “j” to indicate any final state quark, i.e. including also third generation.
Figure 3: Example Feynman diagrams for single and pair production of VLQs and their decays.
Both ATLAS and CMS have searched for such vector-like quarks in a number of publications. For example, ATLAS has searched for pair-produced VLQs decaying to 1st generation quarks in ATLAS:2024zlo . Limits are given as function of the VLQ mass, for different branching ratios Br($Q\to W+j$) versus Br($Q\to H+j$), assuming Br($Q\to W+j$)+Br($Q\to H+j$)+Br($Q\to Z+j$)=1=1 ATLAS:2024zlo . Derived limits are in the range (9001500)(900-1500) GeV for Br($Q\to W+j$) in the range (10-100) %. In ATLAS:2024gyc ATLAS presented the results of a search for pair-produced VLQs decaying to third generation quarks. Limits range from (11.7)(1-1.7) TeV depending mostly on the branching ratio Br($Q\to W+t/b$). These limits are slightly stronger than the 1st generation quark limits, mostly due to lower backgrounds.
There are also searches for singly produced VLQs. However, in the search ATLAS:2022ozf ATLAS assumes that the VLQ decays to the final state $t+h$ or $W+b$, where the top/bottom is tagged in order to reduce backgrounds. Similarly CMS CMS:2024qdd searched for single VLQ production in the final state $t+h$ and $t+Z$. For $\kappa_{T}=1$, where $\kappa_{T}$ measures the strength of the interaction of the VLQ with electro-weak gauge bosons relative to standard EW couplings, a lower limit on VLQs decaying to top quarks of around $m_{\rm VLQ}\geq 1.9$ TeV is found in ATLAS:2022ozf . There are no searches yet published for singly produced VLQs decaying to second or 1st generation quarks. However, due to much larger backgrounds in such searches, one can expects limits would be considerably weaker.
However, we need to stress that all the VLQ searches discussed up to now depend on the presence of mixing. The neutron decay diagrams, that we study here, however, do not require any of the Yukawa coupling in Eq. (24) to be non-zero. Instead, couplings of the form $F$-$q$-$S$, where $F$, $q$ and $S$ stand for a VLQ, a SM quark and one of the BSM scalars must be present. If the VLQs are pair produced at the LHC they can then decay to a final states containing $q+N_{R}+N_{R}$, i.e. jet plus missing energy. This signal is equivalent to the one used in the standard squark search at the LHC. For example, ATLAS ATLAS:2020syg has searched for pair produced squarks with ${\cal L}=139/$fb. ATLAS derives limits for one generation of squarks (roughly $m_{\tilde{q}}\geq 1.2$ TeV) and 8 degenerate squarks (roughly $m_{\tilde{q}}\geq 1.8$ TeV) for light neutralinos. Since VLQs (being fermions) have cross sections at least $\sim$ 4 times larger than a colour triplet scalar of the same mass, this SUSY search ATLAS:2020syg could be reinterpreted to yield a lower limit on the VLQ mass. We estimate this limit to be roughly $m_{\rm VLQ}\geq 1.5$ TeV, for any of the three coloured VLQs under consideration.
The coloured scalar states appear either as leptoquarks or as diquarks in the different decompositions, discussed in the previous sections. The leptoquark couplings are:
${\cal L}^{\rm LQ}=$ $Y_{Nu}\ \overline{{u_{R}}^{c}}N_{R}S^{\dagger}_{3,1,2/3}+Y_{de}\ \overline{{d_{R}}^{c}}e_{R}S_{3,1,2/3}$
$+$ $Y_{Nd}\ \overline{{d_{R}}^{c}}N_{R}S^{\dagger}_{3,1,-1/3}+Y_{eu}\ \overline{{u_{R}}^{c}}e_{R}S^{\dagger}_{3,1,-1/3}+Y_{LQ}\ \overline{Q^{c}}LS^{\dagger}_{3,1,-1/3}+{\rm h.c.},$
while the diquark couplings can be written as:
${\cal L}^{\rm DQ}=Y_{dd}\ \overline{{d_{R}}^{c}}d_{R}S_{3,1,2/3}+Y_{QQ}\ \overline{Q^{c}}QS_{3,1,-1/3}+Y_{ud}\ \overline{{u_{R}}^{c}}d_{R}S_{3,1,-1/3}+{\rm h.c.}$ (26)
Again, we have suppressed generation indices. We stress, however, that the flavour diagonal entries in $Y_{dd}$ are zero, see section III. Importantly, while the LQs and DQs appear with the same SM quantum numbers, they must be different states, otherwise one can generate $d=6$ proton decay tree-level diagrams, which would lead to lower limits on their masses near the GUT scale and render invisible neutron decay completely negligible, as discussed in section III. Note that only ($Y_{Nu}$, $Y_{Nd}$ and $Y_{ud}$) of the eight couplings given in Eqs. (IV) and (26) appear in the decompositions for the neutron decay. This is important, since none of these couplings will generate final states with charged leptons.
[…]
Single or resonant production of LQs requires large Yukawa couplings to give sizeable cross sections. ATLAS searched for resonant LQs ATLAS:2025upm via a lepton-jet signature, with either one or two leptons in the final state. For $Y_{de}=1.0$ LQs with masses below $m_{LQ}=3.4$ TeV are excluded by this search. CMS searched for single LQs in t-channel diagrams CMS:2025iix , leading to di-lepton final states. Both scalar and vector LQs were considered. For scalar [vector] LQs with masses between (15)(1-5) TeV Yukawa couplings larger than (0.31.0)(0.3–1.0) [(0.11.4)(0.1-1.4)] have been excluded.
[…]
Finally, diquarks are stringently constrained from dijet searches, since the production is s-channel enhanced. For example, CMS CMS:2019gwf gives a lower limit on the mass of a colour triplet scalar of $m_{DQ}\simeq 7.5$ TeV assuming a Yukawa coupling to two quarks of electromagnetic strength, $e$. The limits derived by CMS are strictly speaking valid only for a scalar diquark coupling to both, up-type and down-type quarks with the same strength. Thus, they apply only to $S_{3,1,-1/3}$.
The searches discussed above are based on luminosities up to 140/fb. Moderate improvements can be expected from the high-luminosity LHC. However, the future FCC-hh FCC:2018vvp would considerably improve sensitivities. For example, FCC:2018byv quotes sensitivities up to $m_{LQ}\sim 8$ ($15$) TeV for pair (singly) produced leptoquarks, decaying to charged leptons. Many other search channels, discussed above, should improve by similar factors.
We will close this section with a short discussion of the $d=12$ operator given in Eq. (1). In the beginning of this section we have argued that a Wilson coefficient of $c_{W}=1$ is often unrealistic in UV models, thus EFT tends to overestimate the masses of the BSM states. This is certainly true also for a $d=12$ operator. [^7] Note that in the decomposition of Eq. (1) up to seven different Yukawa couplings appear. We have used an automated code for operator decomposition, based on the diagrammatic method and described in Cepedello:2022pyx ; Cepedello:2023yao , to decompose Eq. (1) at tree-level. At such a large dimension, there is a proliferation of models. We count 39 topologies, 150 different diagrams and 12713 model variants.
Despite this large number of possible model variations all model variants can be described with just 38 different fields (scalars and fermions, no vectors). 26 of these are triplets of colour, either diquarks, leptoquarks or heavy vector-like quarks. All scalars and fermions discussed above appear in this list too and the constraints we have discussed above apply also to this case. The remaining coloured states, not covered in the above discussion will have bounds, which are at least as strong, since they are all either larger $SU(2)$ multiplets or states with larger hyper-charge. There is, however, one important difference between the UV states of the $d=12$ operator and those of the d=(79)d=(7-9) operators. Whereas for the $d=9$ operators, strictly speaking, only right-handed neutrino final states are required by the couplings appearing in the decomposition, in the case of the $d=12$ operator similar decay rates of the leptoquarks to final states with missing energy and final states with charged leptons are expected. As discussed above, final states with charged leptons are a better signal than missing energy at the LHC and thus give generally more stringent limits. In other words, limits on states from the decompositions of the $d=12$ operator from LHC will in general be more stringent than those for the $d=9$ case, that motivated the discussion given in this section.
V Summary
In this paper we have studied invisible neutron decay. Limits on nucleon decay modes with three charged leptons Super-Kamiokande:2020tor are much stronger than those of invisible neutron decay ParticleDataGroup:2024cfk . Invisible neutron decay with observable rates can therefore be generated only by operators that do not also generate charged lepton final states. In SMEFT one has to go up to $d=12$ to find such an operator. Current limits on invisible neutron decay correspond then to scales of order $\Lambda\simeq 13$ TeV. An observation of invisible neutron decay in the next round of experiments might then be accompanied by new physics at the future FCC-hh FCC:2018byv .
However, this conclusion changes, if one adds new light degrees of freedom to the SM particle content. We have discussed four different extensions of the SM for which invisible neutron decay with observable rates can be induced by operators of dimension d=(79)d=(7-9). We have estimated the half-lives in each case and for each model we also give all possible UV completions at tree-level. Current limits on the scales of these operators range from $\Lambda\simeq 180$ TeV for the variant model with only right-handed neutrinos ($d=9$ operator) to $\Lambda\simeq 2\cdot 10^{9}$ GeV ($d=7$ operator).
A discovery of invisible neutron decay, therefore may point towards to existence of new, light BSM states. In particular, we stress that in all possible models a neutral fermion state with the quantum numbers of a right-handed neutrino must exist. Thus, these models will also be able to explain neutrino masses.
We have also discussed constraints on the allowed rate for invisible neutron decay, that could be placed by a search for $n\to\pi^{0}+E\hskip-6.00006pt/\;\>$. Hyper-Kamiokande Hyper-Kamiokande:2018ofw should be able to improve on the limit of $n\to\pi^{0}+E\hskip-6.00006pt/\;\>$ from Super-Kamiokande Super-Kamiokande:2013rwg by a considerable factor. We note that the Super-Kamiokande analysis Super-Kamiokande:2013rwg gives only a limit on a two-body final state, but the same type of search can also be used to constrain decays with more than one invisible particle in the final state. Non-observation of $n\to\pi^{0}+3\bar{N}$ at Hyper-Kamiokande might rule out the possibility of observing invisible neutron decay in the foreseeable future.
However, to end this discussion on a more positive note, we mention that the sensitivity of Hyper-Kamiokande and the future JUNO and (the proposed) THEIA experiment to the scales of the invisible neutron decay operators will be rather similar. Observation of invisible neutron decay in JUNO should lead to an excess in the Hyper-Kamiokande search for $n\to\pi^{0}+E\hskip-6.00006pt/\;\>$ and vice versa. The question whether invisible neutron decay or the decay $n\to\pi^{0}+E\hskip-6.00006pt/\;\>$ would be discovered first, however, is not so easy to answer. While the future Hyper-Kamiokande experiment Hyper-Kamiokande:2018ofw is much larger than Super-Kamiokande, and thus allows to probe much larger half-lives, already the Super-Kamiokande limit Super-Kamiokande:2013rwg is background dominated. To claim a discovery in this case requires both hundreds of signal events and a detailed understanding of backgrounds. JUNO’s search for the invisible neutron decay JUNO:2024pur , on the other hand, is based on a characteristic triple coincidence arising from the invisible decays of s-shell neutrons in 12C, which leaves very few background events, making a discovery with just a handful of events possible.
Finally, one can speculate that adding the two gammas from the $\pi^{0}$ decay to the triple coincidence for the invisible neutron decay would eliminate all remaining background in JUNO’s search. We do not know, however, whether this advantage is (over-)compensated by the loss of signal extraction efficiency or not. Only the experimental collaboration can perform an analysis sufficiently sophisticated to decide whether $n\to\pi^{0}+E\hskip-6.00006pt/\;\>$ or the invisible neutron decay itself would be more sensitive to the different models (and the corresponding operators) that we discussed in the present paper.
Acknowledgements
[…]
Appendix A List of the high-energy completions
Here we list the decompositions of the effective operators and the charges under the SM gauge symmetries of the necessary mediator fields in each decomposition.
  • Variant I with light right-handed neutrinos $N_{R}$ — Tabs. 1, 3, 4, and 5. Fig. 2 for the topologies.
  • Variant II with an $N_{R}$ and an axion-like particle $a$ — Tabs. 6 and 7. Fig. 4 for the topology.
  • Variant III with an $N_{R}$ and a scalar field $\phi$ — Tabs. 8 and 9. Fig. 4 for the topology.
  • Variant IV with an $N_{R}$ and a vector boson $Z^{\prime}$ — The mediators appear in this variant are the same as those listed in Tabs. 6 and 7 for Variant II.
The “Decomposition” column shows how the outer fields are distributed to the vertices in the corresponding topology. The fields in a parenthesis are in the same vertex. Let us take the first line of Tab. 3
$(u_{R}d_{R})(N_{Ri})(N_{Rj})(d_{R}N_{Rk})$ (27)
as an example. $(u_{R}d_{R})$” means that $u_{R}$ and $d_{R}$ correspond to the two outer fields in the leftmost vertex in Topology B given in Fig. 2. The mediator field between the $(u_{R}d_{R})$ vertex and the second-left vertex with $N_{Ri}$ is $S_{1}(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{1}}},+1/3)$. The direction of the $S_{1}$ particle number is indicated with arrow in Fig. 2. The two of the middle vertices with $N_{Ri}$ and $N_{Rj}$ are mediated by the fermion $F(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{1}}},+1/3)$. The outer fields in the rightmost vertex are $d_{R}$ and $N_{Rk}$, and the mediator between the $(N_{Rj})$ vertex and the $(d_{R}N_{Rk})$ vertex is $S_{2}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},-1/3)$. In short, the high-energy completion of this decomposition is determined as
[…]
$+y_{\overline{F}NS_{1}}(\overline{F_{L}})_{I\dot{a}}(N_{R})^{\dot{a}}(S_{1})^{I}+y_{NFS_{2}}(\overline{{N_{R}}^{c}})_{\dot{a}}(F_{R})^{I\dot{a}}(S_{2})_{I}.$ (28)
The charges of $S_{2}$ and $S_{1}^{\dagger}$ are the same. However, if they are an identical field, it mediates the $d=6$ effective operator
$\mathscr{L}_{d=6}=\frac{y_{ud}y_{dN}}{M_{S_{1}}^{2}}\epsilon^{IJK}(\overline{{u_{R}}^{c}})_{I\dot{a}}(d_{R})_{J}^{\dot{a}}(\overline{{d_{R}}^{c}})_{K\dot{b}}(N_{R})^{\dot{b}}$ (29)
which induces $p\rightarrow\pi^{+}\bar{N}$ with larger decay rate than $d=9$ induced invisible neutron decay, i.e., the strong bound from $p\rightarrow\pi^{+}\bar{N}$ constrains the invisible neutron decay to an unreachable level. In short, if one wants to have this decomposition for the invisible neutron decay, the mediator $S_{2}$ must be an independent field from $S_{1}^{\dagger}$, although the SM charges are the same. This is mentioned in the “Comment” column as $S_{2}\neq S_{1}^{\dagger}$ to avoid d=6d=6”.
The decompositions with the comment $d_{R}d_{R}S^{\dagger}=0$” contain the interaction
$\mathscr{L}=y\epsilon^{IJK}(\overline{{d_{R}}^{c}})_{I\dot{a}}(d_{R})_{J}^{\dot{a}}(S^{\dagger})_{K}$ (30)
which vanishes because of the colour antisymmetric nature of the two $d_{R}$s forming a scalar. The decomposition $(u_{R}N_{Ri})(d_{R})(d_{R})(N_{Rj}N_{Rk})$ does not contain the interaction Eq. (30), but the effective operator induced from this decomposition results in the same structure $\epsilon^{IJK}(\overline{{d_{R}}^{c}})_{I\dot{a}}(d_{R})_{J}^{\dot{a}}$ that vanishes.
For all fermion mediators that appear in the decompositions, we assume a vector-like fermions (left and right 2-spinors with the same charges form a 4-componenet Dirac fermion). However, the mediator of the SM singlet field can be a Majorana fermion (only 2 components out of 4 are independent). The decompositions with the SM singlet fermion mediator are indicated with the comment “$F$ can be Majorana”.
In the last decomposition in Tab. 4, the two vector mediators have the same SM charges. Since they are antisymmetric under both $SU(3)_{c}$ and $SU(2)_{L}$, the interaction $V_{1}V_{2}S_{3}$
[…]
does not vanish even if they are identical fields. This is different from the last decomposition in Tab. 1, where $S_{3}\neq S_{2}$ is required so that the $S_{1}S_{2}S_{3}$ interaction does not vanish. The comment $V_{1}=V_{2}$” indicates this fact.
Decomposition $S_{1}$ $F$ $S_{2}$ Comment
$(u_{R}d_{R})(N_{Ri})(N_{Rj})(d_{R}N_{Rk})$ $(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{1}}},+1/3)$ $(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{1}}},+1/3)$ $(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},-1/3)$ $S_{2}\neq S_{1}^{\dagger}$ to avoid $d=6$
$(u_{R}d_{R})(d_{R})(N_{Ri})(N_{Rj}N_{Rk})$ $(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{1}}},+1/3)$ $(\boldsymbol{{\mathrm{1}}},\boldsymbol{{\mathrm{1}}},0)$ $(\boldsymbol{{\mathrm{1}}},\boldsymbol{{\mathrm{1}}},0)$ $F$ can be Majorana
$(u_{R}d_{R})(N_{Ri})(d_{R})(N_{Rj}N_{Rk})$ $(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{1}}},+1/3)$ $(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{1}}},+1/3)$ $(\boldsymbol{{\mathrm{1}}},\boldsymbol{{\mathrm{1}}},0)$
$(d_{R}N_{Ri})(u_{R})(d_{R})(N_{Rj}N_{Rk})$ $(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},-1/3)$ $(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{1}}},+1/3)$ $(\boldsymbol{{\mathrm{1}}},\boldsymbol{{\mathrm{1}}},0)$
$(d_{R}N_{Ri})(d_{R})(u_{R})(N_{Rj}N_{Rk})$ $(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},-1/3)$ $(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{1}}},-2/3)$ $(\boldsymbol{{\mathrm{1}}},\boldsymbol{{\mathrm{1}}},0)$
$(u_{R}N_{Ri})(N_{Rj})(N_{Rk})(d_{R}d_{R})$ $(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},+2/3)$ $(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},+2/3)$ $(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{1}}},-2/3)$ $d_{R}d_{R}S_{2}^{\dagger}=0$
$(u_{R}N_{Ri})(d_{R})(d_{R})(N_{Rj}N_{Rk})$ $(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},+2/3)$ $(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{1}}},+1/3)$ $(\boldsymbol{{\mathrm{1}}},\boldsymbol{{\mathrm{1}}},0)$ $\epsilon^{IJK}(d_{RI}d_{RJ})=0$
$(d_{R}d_{R})(u_{R})(N_{Ri})(N_{Rj}N_{Rk})$ $(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{1}}},-2/3)$ $(\boldsymbol{{\mathrm{1}}},\boldsymbol{{\mathrm{1}}},0)$ $(\boldsymbol{{\mathrm{1}}},\boldsymbol{{\mathrm{1}}},0)$ $d_{R}d_{R}S_{1}^{\dagger}=0$
$(d_{R}d_{R})(N_{Ri})(u_{R})(N_{Rj}N_{Rk})$ $(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{1}}},-2/3)$ $(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{1}}},-2/3)$ $(\boldsymbol{{\mathrm{1}}},\boldsymbol{{\mathrm{1}}},0)$ $d_{R}d_{R}S_{1}^{\dagger}=0$
$(u_{R}N_{Ri})(d_{R})(N_{Rj})(d_{R}N_{Rk})$ $(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},+2/3)$ $(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{1}}},+1/3)$ $(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},-1/3)$
$(u_{R}N_{Ri})(N_{Rj})(d_{R})(d_{R}N_{Rk})$ $(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},+2/3)$ $(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},+2/3)$ $(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},-1/3)$
$(d_{R}N_{Ri})(u_{R})(N_{Rj})(d_{R}N_{Rk})$ $(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},-1/3)$ $(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{1}}},+1/3)$ $(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},-1/3)$
Table 3: Decompositions and the necessary mediator fields of the $d=9$ effective operator $u_{R}d_{R}d_{R}N_{R}N_{R}N_{R}$ with Topology B in Fig. 2.
Decomposition $S_{1}/V_{1}$ $S_{2}/V_{2}$ $S_{3}/V_{3}$ Comment
$(QQ)(d_{R}N_{Ri})(N_{Rj}N_{Rk})$ $S_{1}(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{1}}},+1/3)$ $S_{2}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},-1/3)$ $S_{3}(\boldsymbol{{\mathrm{1}}},\boldsymbol{{\mathrm{1}}},0)$
$(Qd_{R})(QN_{Ri})(N_{Rj}N_{Rk})$ $V_{1}(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{2}}},-1/6)$ $V_{2}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{2}}},+1/6)$ $S_{3}(\boldsymbol{{\mathrm{1}}},\boldsymbol{{\mathrm{1}}},0)$
$(QN_{Ri})(QN_{Rj})(d_{R}N_{Rk})$ $V_{1}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{2}}},+1/6)$ $V_{2}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{2}}},+1/6)$ $S_{3}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},-1/3)$ $V_{1}=V_{2}$
Table 4: Decompositions of the $QQd_{R}N_{R}N_{R}N_{R}$ operator with Topology A in Fig. 2.
Decomposition $S_{1}/V_{1}$ $F$ $S_{2}/V_{2}$ Comment
$(QQ)(N_{Ri})(N_{Rj})(d_{R}N_{Rk})$ $S_{1}(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{1}}},+1/3)$ $(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{1}}},+1/3)$ $S_{2}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},-1/3)$
$(QQ)(d_{R})(N_{Ri})(N_{Rj}N_{Rk})$ $S_{1}(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{1}}},+1/3)$ $(\boldsymbol{{\mathrm{1}}},\boldsymbol{{\mathrm{1}}},0)$ $S_{2}(\boldsymbol{{\mathrm{1}}},\boldsymbol{{\mathrm{1}}},0)$ $F$ can be Majorana
$(QQ)(N_{Ri})(d_{R})(N_{Rj}N_{Rk})$ $S_{1}(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{1}}},+1/3)$ $(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{1}}},+1/3)$ $S_{2}(\boldsymbol{{\mathrm{1}}},\boldsymbol{{\mathrm{1}}},0)$
$(d_{R}N_{Ri})(Q)(Q)(N_{Rj}N_{Rk})$ $S_{1}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},-1/3)$ $(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{2}}},-1/6)$ $S_{2}(\boldsymbol{{\mathrm{1}}},\boldsymbol{{\mathrm{1}}},0)$
$(Qd_{R})(N_{Ri})(N_{Rj})(QN_{Rk})$ $V_{1}(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{2}}},-1/6)$ $(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{2}}},-1/6)$ $V_{2}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{2}}},+1/6)$
$(Qd_{R})(Q)(N_{Ri})(N_{Rj}N_{Rk})$ $V_{1}(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{2}}},-1/6)$ $(\boldsymbol{{\mathrm{1}}},\boldsymbol{{\mathrm{1}}},0)$ $S_{2}(\boldsymbol{{\mathrm{1}}},\boldsymbol{{\mathrm{1}}},0)$ $F$ can be Majorana
$(Qd_{R})(N_{Ri})(Q)(N_{Rj}N_{Rk})$ $V_{1}(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{2}}},-1/6)$ $(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{2}}},-1/6)$ $S_{2}(\boldsymbol{{\mathrm{1}}},\boldsymbol{{\mathrm{1}}},0)$
$(QN_{Ri})(Q)(d_{R})(N_{Rj}N_{Rk})$ $V_{1}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{2}}},+1/6)$ $(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{1}}},+1/3)$ $S_{2}(\boldsymbol{{\mathrm{1}}},\boldsymbol{{\mathrm{1}}},0)$
$(QN_{Ri})(d_{R})(Q)(N_{Rj}N_{Rk})$ $V_{1}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{2}}},+1/6)$ $(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{2}}},-1/6)$ $S_{2}(\boldsymbol{{\mathrm{1}}},\boldsymbol{{\mathrm{1}}},0)$
$(QN_{Ri})(d_{R})(N_{Rj})(QN_{Rk})$ $V_{1}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{2}}},+1/6)$ $(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{2}}},-1/6)$ $V_{2}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{2}}},+1/6)$ $V_{1}=V_{2}$
$(QN_{Ri})(Q)(N_{Rj})(d_{R}N_{Rk})$ $V_{1}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{2}}},+1/6)$ $(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{1}}},+1/3)$ $S_{2}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},-1/3)$
$(QN_{Ri})(N_{Rj})(Q)(d_{R}N_{Rk})$ $V_{1}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{2}}},+1/6)$ $(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{2}}},+1/6)$ $S_{2}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},-1/3)$
Table 5: Decompositions of the $QQd_{R}N_{R}N_{R}N_{R}$ operator with Topology B in Fig. 2.
Figure 4: Topology for the operators with four fermions and 1 scalar, such as $u_{R}d_{R}d_{R}N_{R}\phi$ and $u_{R}d_{R}d_{R}\overline{N_{R}}a$, i.e., One of the outer legs should be a scalar, $\phi$ or $a$. The directions of arrows on the outer legs depends on the decomposition. Although the mediators are given with solid lines in this topology, they can be scalars $S$, vectors $V$ or fermions $F$, depending on the distribution of outer fields.
Decomposition $S_{1}/V_{1}/F_{1}$ $S_{2}/V_{2}/F_{2}$ Comment
$\partial](d_{R}d_{R})(u_{R})(\overline{N_{R}}a)$ $S_{1}(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{1}}},-2/3)$ $F_{2}(\boldsymbol{{\mathrm{1}}},\boldsymbol{{\mathrm{1}}},0)$ $d_{R}d_{R}S_{1}^{\dagger}=0$
$\partial](d_{R}\overline{N_{R}})(u_{R})(d_{R}a)$ $V_{1}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},-1/3)$ $F_{2}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},-1/3)$
$\partial](u_{R}d_{R})(d_{R})(\overline{N_{R}}a)$ $S_{1}(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{1}}},+1/3)$ $F_{2}(\boldsymbol{{\mathrm{1}}},\boldsymbol{{\mathrm{1}}},0)$ $F_{2}$ can be Majorana
$\partial](u_{R}\overline{N_{R}})(d_{R})(d_{R}a)$ $V_{1}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},+2/3)$ $F_{2}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},-1/3)$
$\partial](u_{R}a)(d_{R})(d_{R}\overline{N_{R}})$ $F_{1}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},+2/3)$ $V_{2}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},-1/3)$
$\partial](u_{R}d_{R})(\overline{N_{R}})(d_{R}a)$ $S_{1}(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{1}}},+1/3)$ $F_{2}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},-1/3)$
$\partial](u_{R}a)(\overline{N_{R}})(d_{R}d_{R})$ $F_{1}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},+2/3)$ $S_{2}(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{1}}},-2/3)$ $d_{R}d_{R}S_{2}^{\dagger}=0$
$\partial](u_{R}d_{R})(a)(d_{R}\overline{N_{R}})$ $S_{1}(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{1}}},+1/3)$ $V_{2}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},-1/3)$
$\partial](u_{R}\overline{N_{R}})(a)(d_{R}d_{R})$ $V_{1}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},+2/3)$ $S_{2}(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{1}}},-2/3)$ $d_{R}d_{R}S_{2}^{\dagger}=0$
Table 6: Decompositions of $u_{R}d_{R}d_{R}\overline{N_{R}}a$, which result in the $d=8$ effective operator with a derivative operator, $\partial au_{R}d_{R}d_{R}\overline{N_{R}}$. The operator $u_{R}d_{R}d_{R}\overline{N_{R}}Z^{\prime}$ with a vector boson is decomposed with the same mediators.
Decomposition $S_{1}/V_{1}/F_{1}$ $S_{2}/V_{2}/F_{2}$ Comment
$\partial](QQ)(d_{R})(\overline{N_{R}}a)$ $S_{1}(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{1}}},+1/3)$ $_{2}(\boldsymbol{{\mathrm{1}}},\boldsymbol{{\mathrm{1}}},0) FF$ can be Majorana
$\partial](Q\overline{N_{R}})(d_{R})(Qa)$ $S_{1}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{2}}},+1/6)$ $F_{2}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{2}}},+1/6)$
$\partial](Qd_{R})(Q)(\overline{N_{R}}a)$ $V_{1}(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{2}}},-1/6)$ $_{2}(\boldsymbol{{\mathrm{1}}},\boldsymbol{{\mathrm{1}}},0) FF$ can be Majorana
$\partial](Q\overline{N_{R}})(Q)(d_{R}a)$ $S_{1}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{2}}},+1/6)$ $F_{2}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},-1/3)$
$\partial](Qa)(Q)(d_{R}\overline{N_{R}})$ $F_{1}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{2}}},+1/6)$ $V_{2}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},-1/3)$
$\partial](Qd_{R})(\overline{N_{R}})(Qa)$ $V_{1}(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{2}}},-1/6)$ $F_{2}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{2}}},+1/6)$
$\partial](d_{R}a)(\overline{N_{R}})(QQ)$ $F_{1}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},-1/3)$ $S_{2}(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{1}}},+1/3)$
$\partial](Qd_{R})(a)(Q\overline{N_{R}})$ $V_{1}(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{2}}},-1/6)$ $S_{2}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{2}}},+1/6)$
$\partial](d_{R}\overline{N_{R}})(a)(QQ)$ $V_{1}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},-1/3)$ $S_{2}(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{1}}},+1/3)$
Table 7: Decompositions of $QQd_{R}\overline{N_{R}}a$, which result in the $d=8$ effective operator with a derivative operator, $\partial aQQd_{R}\overline{N_{R}}$. The operator $QQd_{R}\overline{N_{R}}Z^{\prime}$ is decomposed with the same mediators.
Decomposition $S_{1}/F_{1}$ $S_{2}/F_{2}$ Comment
$(d_{R}d_{R})(u_{R})(N_{R}\phi)$ $S_{1}(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{1}}},-2/3)$ $F_{2}(\boldsymbol{{\mathrm{1}}},\boldsymbol{{\mathrm{1}}},0)$ $d_{R}d_{R}S_{1}^{\dagger}=0$
$(d_{R}N_{R})(u_{R})(d_{R}\phi)$ $S_{1}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},-1/3)$ $F_{2}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},-1/3)$
$(u_{R}d_{R})(d_{R})(N_{R}\phi)$ $S_{1}(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{1}}},+1/3)$ $F_{2}(\boldsymbol{{\mathrm{1}}},\boldsymbol{{\mathrm{1}}},0)$ $F_{2}$ can be Majorana
$(u_{R}N_{R})(d_{R})(d_{R}\phi)$ $S_{1}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},+2/3)$ $F_{2}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},-1/3)$
$(u_{R}\phi)(d_{R})(d_{R}N_{R})$ $F_{1}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},+2/3)$ $S_{2}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},-1/3)$
$(u_{R}d_{R})(N_{R})(d_{R}\phi)$ $S_{1}(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{1}}},+1/3)$ $F_{2}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},-1/3)$
$(u_{R}\phi)(N_{R})(d_{R}d_{R})$ $F_{1}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},+2/3)$ $S_{2}(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{1}}},-2/3)$ $d_{R}d_{R}S_{2}^{\dagger}=0$
$(u_{R}d_{R})(\phi)(d_{R}N_{R})$ $S_{1}(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{1}}},+1/3)$ $S_{2}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},-1/3)$ $S_{2}^{\dagger}\neq S_{1}$ to avoid $d=6$ $u_{R}d_{R}d_{R}N_{R}$
$(u_{R}N_{R})(\phi)(d_{R}d_{R})$ $S_{1}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},+2/3)$ $S_{2}(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{1}}},-2/3)$ $d_{R}d_{R}S_{2}^{\dagger}=0$
Table 8: Decompositions of $u_{R}d_{R}d_{R}N_{R}\phi$.
Decomposition $S_{1}/V_{1}/F_{1}$ $S_{2}/V_{2}/F_{2}$ Comment
$(Qd_{R})(Q)(N_{R}\phi)$ $V_{1}(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{2}}},-1/6)$ $F_{2}(\boldsymbol{{\mathrm{1}}},\boldsymbol{{\mathrm{1}}},0)$ $F_{2}$ can be Majorana
$(QN_{R})(Q)(d_{R}\phi)$ $V_{1}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{2}}},+1/6)$ $F_{2}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},-1/3)$
$(Q\phi)(Q)(d_{R}N_{R})$ $F_{1}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{2}}},+1/6)$ $S_{2}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},-1/3)$
$(QQ)(d_{R})(N_{R}\phi)$ $S_{1}(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{1}}},+1/3)$ $F_{2}(\boldsymbol{{\mathrm{1}}},\boldsymbol{{\mathrm{1}}},0)$ $F_{2}$ can be Majorana
$(QN_{R})(d_{R})(Q\phi)$ $V_{1}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{2}}},+1/6)$ $F_{2}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{2}}},+1/6)$
$(QQ)(N_{R})(d_{R}\phi)$ $S_{1}(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{1}}},+1/3)$ $F_{2}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},-1/3)$
$(Qd_{R})(N_{R})(Q\phi)$ $V_{1}(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{2}}},-1/6)$ $F_{2}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{2}}},+1/6)$
$(QQ)(\phi)(d_{R}N_{R})$ $S_{1}(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{1}}},+1/3)$ $S_{2}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{1}}},-1/3)$ $S_{2}^{\dagger}\neq S_{1}$ to avoid $d=6$ $QQd_{R}N_{R}$
$(Qd_{R})(\phi)(QN_{R})$ $V_{1}(\overline{\boldsymbol{{\mathrm{3}}}},\boldsymbol{{\mathrm{2}}},-1/6)$ $V_{2}(\boldsymbol{{\mathrm{3}}},\boldsymbol{{\mathrm{2}}},+1/6)$ $V_{2}^{\dagger}\neq V_{1}$ to avoid $d=6$ $QQd_{R}N_{R}$
Table 9: Decompositions of $QQd_{R}N_{R}\phi$.